# Simulating Prosocial Behavior and Social Contagion in LLM Agents under Institutional Interventions

### v1

@@ Abstract
As large language models (LLMs) increasingly serve as autonomous agents in social contexts, understanding their capacity for prosocial behavior becomes essential. We present ProSim, a simulation framework designed to examine how prosocial behavior emerges, adapts, and erodes in LLM-based agents under diverse social and institutional conditions. The framework comprises four components: individual simulation, scenario simulation, interaction simulation, and intervention simulation. We conduct three progressive studies to evaluate prosocial alignment. First, we show that LLM agents can demonstrate stable and context-sensitive prosocial behavior across diverse scenarios and adapt their responses under normative policy interventions. Second, we find that agents engage in fairness-based third-party punishment and respond systematically to variations in inequity magnitude and enforcement cost. Third, we show that policy-induced inequities suppress prosocial behavior, propagate through social networks, and are mediated by agents’ perceptions of unfairness. These findings lay the groundwork for evaluating social alignment and modeling institutional dynamics in agent-driven societies.
1 Introduction
@@
Large language models (LLMs) have evolved beyond next-word prediction into general-purpose agents equipped with complex reasoning, decision-making, and social interaction capabilities zhao2023survey ; gpt4 . Recent studies suggest that LLMs can exhibit emergent social-cognitive abilities mou2024individual ; piao2025agentsociety , including theory of mind li2023theory ; strachan2024testing , moral judgment ramezani2023knowledge , and value alignment liu2022aligning . These developments position LLMs as promising tools for simulating human-like behavior in synthetic populations.
This work focuses on prosocial behavior, defined as voluntary actions intended to benefit others or promote collective welfare penner2005prosocial ; grueneisen2022development . Prosocial behavior plays a critical role in sustaining cooperation, trust, and social cohesion cameron2022empathy ; carlo2005interplay ; thielmann2020personality . Prior research brookins2023playing ; xie2024can has primarily examined prosociality in LLMs through economic game paradigms such as the dictator game and public goods game. However, existing studies face two key limitations. First, they narrowly frame prosociality in terms of economic cooperation (e.g., resource allocation), overlooking diverse real-world expressions such as donating and volunteering. Second, they rely on static, single-shot evaluations, neglecting how prosocial behavior dynamically evolves in changing environments or in response to institutional interventions.
@@
To address these limitations, we propose ProSim, a comprehensive simulation framework that models the emergence and evolution of prosocial behavior in LLM-based agents. The framework consists of four components. (1) Individual simulation module, which instantiates each agent with demographic attributes and psychological traits such as empathy and moral identity. (2) interaction simulation module, which places agents in a small-world network and supports repeated multi-agent interactions. (3) Scenario simulation module, which reproduces six distinct prosocial tasks: helping, donating, volunteering, cooperating, information sharing, and recycling. (4) Intervention simulation module, which implements prosocial policy interventions and allows manipulation of fairness conditions through reward asymmetry and burden asymmetry. Building on this framework, we conduct a series of simulation studies to investigate the social capacities of LLM agents from multiple perspectives.
@@
We begin by evaluating whether LLM agents naturally exhibit prosocial behavior in structured contexts and how they adjust when exposed to policy-based interventions. Drawing from established behavioral paradigms, we design six typical prosocial scenarios and observe the agents’ default tendencies. To assess their responsiveness to external cues, we introduce four types of prosocial policy interventions and evaluate how these modulate agent behavior. Our results show that LLMs agents can exhibit stable and context-sensitive prosocial behavior across diverse scenarios, and adjust their behavior in response to policy interventions.
@@
In the second study, we test whether LLM agents can perceive and respond to inequity by enforcing social norms. We adapt a third-party punishment paradigm fehr2004third in which agents observe unfair resource distributions and must decide whether to penalize the transgressor at a personal cost. In addition to behavioral responses, we analyze agents’ emotional expressions to assess their affective alignment with human fairness reasoning. Using a comparable human dataset for benchmarking, we find that LLM agents are capable of norm-enforcing third-party punishment, showing sensitivity to both the degree of unfairness and the cost of enforcement.
@@
Finally, we explore how prosocial behavior evolves over time under policy-induced inequity in networked environments. Agents are embedded in a small-world social network and repeatedly interact across rounds of simulated exchanges. Two types of structural unfairness (reward asymmetry and burden asymmetry) are introduced and allowed to diffuse through the network. We track the resulting behavioral trajectories to assess whether prosociality is sustained or eroded at the population level. The findings reveal that Policy inequities significantly undermine prosocial behavior in LLM agents, an effect amplified through social contagion and mediated by agents’ perceived unfairness.
Together, these findings suggest that LLM agents can serve as scalable proxies for studying the emergence and evolution of prosocial behavior in synthetic societies. Our contributions are threefold:
@@
We propose ProSim, a modular simulation framework for modeling the emergence and evolution of prosocial behavior in LLM agents. The framework integrates four key modules to approximate the complexity of real human social environments.
We show that LLM agents not only exhibit prosocial behavior but also adapt their responses based on fairness cues, aligning with human behavioral benchmarks.
Leveraging the aligned simulation environment, we model how structural policy inequities affect the emergence, decay, and diffusion of prosocial norms across simulated societies.
2 The ProSim Framework
Simulating prosocial behavior poses a fundamental challenge due to its psychological and social complexity. Human prosociality is shaped not only by internal traits such as empathy and moral identity, but also by contextual factors, social influence, and institutional structures.
@@ 2 The ProSim Framework
To address this, we introduce ProSim, a modular simulation framework for studying how prosocial behavior emerges, adapts, and deteriorates in LLM-based agents across varied social and policy environments. Grounded in computational social psychology edelmann2020computational , ProSim supports controlled experimentation at both individual and collective levels. As shown in Figure 1, the framework comprises four modules: (1) Individual Simulation, (2) Scenario Simulation, (3) Interaction Simulation, and (4) Intervention Simulation. These components jointly model the dynamic interplay between psychological traits, situational context, and institutional interventions in shaping agent behavior.
2.1 Individual Simulation
@@
This module defines the foundational identity of each agent by capturing the heterogeneity inherent in human populations. Agents are initialized with realistic demographic and psychological profiles, which serve as stable priors for downstream behavior.
@@
Psychological Traits. Each agent is further characterized by two sets of psychological traits: (1) core prosocial dispositions eisenberg1999consistency , including empathic concern, moral identity, altruistic tendency, and social responsibility; and (2) the Big Five personality gosling2003very , including openness, conscientiousness, extraversion, agreeableness, and neuroticism. Trait values are drawn from Gaussian distributions calibrated using meta-analytic norms to reflect empirical variability.
@@
All attributes are encoded into the agent’s natural language prompt, grounding each agent in a contextually rich identity. This design ensures that behavior arises not from static rules but from socially and psychologically plausible profiles. More details are documented in Appendix C.1.
2.2 Scenario Simulation
@@
This module defines the situational contexts in which agents are prompted to make prosocial decisions. This component is essential for evaluating how different types of social dilemmas and environmental cues influence prosocial tendencies under controlled conditions. To ensure comprehensive coverage of real-world prosocial behaviors, we define six distinct scenarios that vary across cost level, norm expectation, reward visibility, and dependency level on others.
Dimension | Volunteering | Helping | Donating | Cooperation | Recycling | Sharing |
Cost Level | High | Medium | Medium | Medium | Low | High |
Norm Expectation | High | High | Medium | Medium | High | Medium |
Reward Visibility | Medium | High | High | High | Low | Medium |
Dependency Level | Medium | High | Low | High | Low | Medium |
As shown in Table 1, the six scenarios (e.g. Volunteering, Helping, Donating, Cooperation, Recycling, Sharing) differ across structural dimensions. The classification of each scenario along the four dimensions is guided by expert judgment from social psychologists. Each scenario is delivered via a standardized prompt, designed to reflect realistic decision-making contexts. Following the prompt, agents respond using a consistent 7-point Likert scale to indicate their level of prosocial intention.
2.3 Interaction Simulation
@@
This module captures how prosocial behavior evolves through repeated agent-to-agent interactions within a structured social network. This layer is essential for modeling emergent dynamics that cannot be explained by individual traits or isolated decisions alone. Concretely, we initialize the agent society using a small-world network generated via the Watts–Strogatz algorithm kleinberg2000small . This structure reflects key properties of real-world social systems, including high local clustering and short average path lengths. Each node represents an LLM-based agent, and each edge denotes a potential communication or observation channel between two agents.
@@
At each simulation timestep , a random subset of edges is activated. Each agent observes the latest prosocial decisions made by neighbors , where is the set of agents connected to via active edges at time . By integrating (i) the shared scenario narrative , (ii) its own prior decision , and (iii) the observed actions of its neighbors , the LLM agent then generates its decision on six prosocial scenarios: Through this module, we simulate how prosocial norms and behaviors diffuse or decay over time within a socially structured community of LLM agents.
2.4 Intervention Simulation
@@
This module enables controlled manipulation of policy conditions to investigate how institutional interventions and structural inequalities influence the development of prosocial behavior. While individual traits and social dynamics contribute to behavioral variation, policy environments play a central role in shaping collective behaviors through top-down regulation.
@@
Policy inequity. Beyond evaluating idealized interventions, we introduce policy inequity to examine the broader social impact of unfair treatment. In real-world systems, rewards and burdens are often distributed asymmetrically across individuals or groups due to structural bias or implementation constraints. To simulate these conditions, we define two types of inequity:
Reward Asymmetry: agents receive unequal recognition or benefits despite contributing equally.
Burden Asymmetry: agents face unequal costs or effort for performing the same prosocial task.
We randomly assign these asymmetries to a subset of agents and allow their effects to propagate through the social network. This setup enables us to assess how inequitable policies influence the emergence and spread of prosocial behavior within agent societies.
2.5 Experimental Settings
@@
To evaluate the capabilities of the ProSim framework, we conduct three progressive studies that examine how LLM agents exhibit, perceive, and respond to prosocial dynamics.
Study 1: Can LLM Agents Exhibit Prosocial Behavior? This study investigates whether LLM agents can demonstrate consistent and context-sensitive prosocial intentions across six scenarios.
Study 2: Do LLM Agents Perceive and Respond to Inequity? This study tests whether agents can detect unequal treatment and adjust their decisions based on the perceived severity of the inequity.
Study 3: How Does Policy Inequity Affect Prosociality? This study explores how policy-induced inequities influence prosocial motivation and how such effects spread through social networks.
@@
Overall Model and Agent Configuration. We evaluate ProSim using five LLMs, including three open-source models (LLaMA-3-8B llama3 , Qwen-2.5-7B qwen2 , DeepSeek-v3 deepseek ) and two proprietary models (GPT-3.5-turbo gpt35 , GPT-4o gpt4 ). LLaMA and Qwen models are deployed on a single NVIDIA A100 GPU for inference, while DeepSeek and the GPT series are accessed via APIs, with the generation temperature is set to 1.0. We initialize 104 agents sampled from real-world distributions, embedded in a small-world network with neighborhood size and rewiring probability . Additional implementation details are provided in each study and Appendix C.
@@
Human Benchmarking. To contextualize LLM behavior, we conduct parallel experiments with 54 human participants recruited online for Study 1 and Study 2. Each participant completed the same tasks presented to the LLM agents, using identical scenario prompts and response formats. Prior to task completion, participants provided informed consent and completed standardized psychological inventories aligned with the trait dimensions used in agent initialization.
3 Can LLM Agents Exhibit Prosocial Behavior?
This section evaluates whether LLM agents can express context-sensitive prosocial behavior in a manner comparable to humans. We first assess their baseline prosocial intentions across diverse scenarios, then investigate how individual psychological traits shape these behaviors. Finally, we examine whether LLMs adapt their responses under explicit prosocial policy interventions.
3.1 Baseline Prosocial Intention in Diverse Scenarios
We begin by assessing the baseline prosocial tendencies of LLM agents across six representative scenarios. Each agent receives a neutral prompt and rates its willingness to perform a prosocial act on a 7-point Likert scale. Figure 2 (left) shows the distribution of prosocial intentions across scenarios. GPT-3.5 exhibits the highest overall prosociality, particularly in tasks involving helping, donating, and cooperation. GPT-4o and DeepSeek-v3 display relative better alignment with human responses across multiple contexts. Qwen-2.5 shows more polarized behavior, responding positively in low-cost or public-good scenarios, but less so in interpersonal settings such as sharing. LLaMA-3 consistently produces the most conservative ratings across all conditions. We also observe that LLMs tend to be more prosocial in scenarios with high normative expectations and less prosocial when the cost of action is high, suggesting a sensitivity to structural dimensions of prosocial tasks. Figure 2 (right) summarizes average prosociality across all tasks. GPT-3.5 ranks highest overall, followed by human participants, with GPT-4o and Qwen-2.5 at comparable levels. DeepSeek-v3 and LLaMA-3 fall lower on the scale. These results suggest that advanced LLMs can express prosocial intentions that approach human-level patterns, although their responses remain sensitive to scenario context and model-specific characteristics. This baseline evaluation sets the stage for further analysis of individual traits and institutional influences on prosocial behavior.
3.2 The Influence of Psychological Traits on Prosociality
@@
3.3 Behavioral Shifts Under Prosocial Policy Interventions
We next evaluate whether LLM agents adjust their prosocial behavior when exposed to external policy cues. Each agent completes the same six scenarios from the baseline condition, with prompts modified to include one of four intervention framings: Economic Incentives, Moral Indoctrination, Regulatory Enforcement, or Social Comparison. We compute the relative change in prosocial intention to assess policy responsiveness. Figure 4 summarizes the results. Intervention effectiveness varies across policy types and models. Regulatory Enforcement yields the most consistent improvement across models and scenarios, suggesting that explicit institutional mandates are broadly effective at eliciting compliance. In contrast, Economic Incentives, while highly effective for human participants, produce mixed effects in LLMs. Only GPT-4o and DeepSeek-v3 exhibit reliable prosocial gains across all six scenarios, though the magnitude of their responses remains below human levels, indicating that monetary cues are less salient for LLMs. At the scenario level, interventions are least effective in the Recycling context. In several models, policy framings even lead to marginal declines in prosociality. This may be due to high baseline prosocial intent in this low-cost, norm-driven scenario, where additional interventions could trigger reactance or reduce intrinsic motivation. The right panel of Figure 4 aggregates the number of positive shifts across all policy-scenario combinations. GPT-4o and DeepSeek-v3 exhibit the strongest and most consistent responsiveness, closely matching human participants. GPT-3.5 and LLaMA-3 show moderate adaptability, while Qwen-2.5 lags behind with the fewest positive responses. Taken together, these findings suggest that advanced LLMs exhibit context-sensitive adjustments to social and normative inputs, a capacity that aligns them more closely with human-like social reasoning.
4 Do LLM Agents Perceive and Respond to Inequity?
This section examines whether LLM agents are capable of recognizing unfairness and responding in norm-enforcing ways. To evaluate this capacity, we adapt a third-party punishment paradigm from behavioral economics fehr2004third , a well-established framework for studying fairness judgment and altruistic punishment. Each agent completes a 60-trial task in which two newly assigned virtual players propose an allocation of $30. Player 1 offers $x to Player 2 and retains $[30-x] for themselves. The participants, acting as third-party judges, then chose between:
@@ 4 Do LLM Agents Perceive and Respond to Inequity?
We record agents’ binary choices (accept = 0, punish = 1) across trials and and compare the results with humans. Emotional responses are also collected but reported separately in Appendix D.
4.1 Third-Party Punishment as a Response to Inequity
We first examine whether LLM agents engage in third-party punishment by rejecting unfair outcomes at a personal cost. Figure 5 shows punishment rates across five LLMs and human participants. DeepSeek-v3 exhibits the highest and most consistent punitive behavior, frequently rejecting unequal allocations regardless of their structure. GPT-4o, GPT-3.5, and LLaMA-3 also show elevated punishment rates, albeit with greater intra-model variability. Qwen-2.5, in contrast, rarely punishes, indicating limited sensitivity to perceived unfairness. Human participants display substantial individual variability, with average punishment levels closely aligning with those of the GPT models. These results suggest that several LLMs can engage in fairness-driven norm enforcement, approximating human behavior in inequity-sensitive contexts. We next assess whether such punishment behavior reflects graded sensitivity to the magnitude of unfairness and associated costs.
4.2 Sensitivity to Inequity Magnitude and Personal Cost
@@
To probe whether LLM agents modulate their punitive behavior based on contextual factors, we manipulate both the allocation scheme and the cost of punishment. Figure 6 illustrates punishment rates across different combinations of fairness levels and punishment costs for each model. Human participants exhibit a clear normative pattern: punishment rates decrease as allocations become more equitable or as costs increase. This behavior reflects a trade-off between norm enforcement and self-interest, consistent with findings from economic game theory. Among LLMs, the GPT series shows the strongest human alignment. These models reliably punish under unfair conditions and abstain under fair ones, demonstrating a consistent threshold-based fairness judgment. Notably, they even exceed human punishment rates in highly unfair trials. DeepSeek-v3 maintains high punishment rates across all unfair conditions, largely ignoring cost variation, suggesting a rigid but robust norm-enforcing strategy. LLaMA-3 also punishes inequity frequently, but its responses are less differentiated by fairness level or cost. Qwen-2.5, meanwhile, shows low overall punishment and lacks sensitivity to either factor. These findings indicate that while some advanced LLMs demonstrate fairness-based reasoning and partial cost sensitivity, others lack the granularity and flexibility observed in human social decision-making.
5 How Does Policy Inequity Affect Prosociality?
Prior analyses show that LLM agents can exhibit prosocial behavior and recognize unfairness. However, it remains unclear how they respond when placed in structurally inequitable environments. This section examines whether policy-induced inequities undermine prosocial motivation, and whether their effects spread through social networks to erode collective prosocial norms.
5.1 Impact of Structural Inequity on Individual Prosocial Behavior

### v2

# Investigating Prosocial Behavior Theory in LLM Agents under Policy-Induced Inequities

@@ Abstract
As large language models (LLMs) increasingly operate as autonomous agents in social contexts, evaluating their capacity for prosocial behavior is both theoretically and practically critical. However, existing research has primarily relied on static, economically framed paradigms, lacking models that capture the dynamic evolution of prosociality and its sensitivity to structural inequities. To address these gaps, we introduce ProSim111Code available at: https://github.com/halsayxi/ProSim/, a simulation framework for modeling the prosocial behavior in LLM agents across diverse social conditions. We conduct three progressive studies to assess prosocial alignment. First, we demonstrate that LLM agents can exhibit human-like prosocial behavior across a broad range of real-world scenarios and adapt to normative policy interventions. Second, we find that agents engage in fairness-based third-party punishment and respond systematically to variations in inequity magnitude and enforcement cost. Third, we show that policy-induced inequities suppress prosocial behavior, propagate norm erosion through social networks. These findings advance prosocial behavior theory by elucidating how institutional dynamics shape the emergence, decay, and diffusion of prosocial norms in agent-driven societies.
Introduction
@@
Large language models (LLMs) have evolved beyond next-word prediction into general-purpose agents equipped with complex reasoning, decision-making, and social interaction capabilities (zhao2023survey; gpt4). Recent studies suggest that LLMs can exhibit emergent social-cognitive abilities (piao2025agentsociety), including theory of mind (strachan2024testing), moral judgment (ramezani2023knowledge), and value alignment (liu2022aligning). These developments position LLMs as promising tools for simulating human-like behavior in synthetic populations.
This work focuses on prosocial behavior theory, defined as voluntary actions intended to benefit others or promote collective welfare (penner2005prosocial). Prosocial behavior is foundational to cooperation, trust, and social cohesion (cameron2022empathy; thielmann2020personality), particularly in addressing societal challenges that demand collective effort. As large language models (LLMs) are increasingly deployed as social agents, evaluating their capacity for prosociality has become both theoretically significant and practically urgent. Prior studies (piatti2024cooperate; xie2024can) have primarily assessed LLM cooperation through economic games such as the dictator game and public goods game. However, this line of work faces two critical limitations. First, it frames prosociality narrowly in terms of economic cooperation, overlooking diverse real-world behaviors that reflect everyday moral engagement. Second, it relies on static, one-shot scenarios, failing to capture how prosocial behavior unfolds over time or adapts to evolving environments, especially under structural inequities such as unfair policy enforcement or social exclusion. These limitations restrict both the theoretical scope and practical utility of current evaluations. To address this gap, we propose a broader and more dynamic framework that examines whether, how, and under what conditions LLM agents can exhibit prosocial behavior across diverse real-world contexts.
@@
In this paper, we propose ProSim, a comprehensive simulation framework that models the emergence and evolution of prosocial behavior in LLM-based agents. The framework consists of four components. (1) Individual simulation module, which instantiates each agent with demographic attributes and psychological traits such as empathy and moral identity. (2) Interaction simulation module, which places agents in a small-world network and supports repeated multi-agent interactions. (3) Scenario simulation module, which reproduces six distinct prosocial tasks: helping, donating, volunteering, cooperating, information sharing, and recycling. (4) Intervention simulation module, which implements prosocial policy interventions and allows manipulation of fairness conditions through reward asymmetry and burden asymmetry. Building on this framework, we conduct a series of simulation studies to investigate the social capacities of LLM agents from multiple perspectives.
@@
We begin by evaluating whether LLM agents naturally exhibit prosocial behavior in structured contexts and how they adjust when exposed to policy-based interventions. Drawing from established behavioral paradigms, we design six typical prosocial scenarios and observe the agents’ default tendencies. To assess their responsiveness to external cues, we introduce four types of prosocial policy interventions and evaluate how these modulate agent behavior. Our results show that LLM agents can exhibit human-like prosocial behavior across diverse scenarios, and adjust their behavior in response to policy interventions.
@@
In the second study, we test whether LLM agents can perceive and respond to inequity by enforcing social norms. We adapt a third-party punishment paradigm (fehr2004third) in which agents observe unfair resource distributions and must decide whether to penalize the transgressor at a personal cost. In addition to behavioral responses, we analyze agents’ emotional expressions to assess their affective alignment with human fairness reasoning. Using a comparable human dataset for benchmarking, we find that LLM agents are capable of norm-enforcing third-party punishment, showing sensitivity to both the degree of unfairness and the cost of enforcement.
@@
Finally, we explore how prosocial behavior evolves over time under policy-induced inequity in networked environments. Agents are embedded in a small-world social network and repeatedly interact across rounds of simulated exchanges. Two types of structural unfairness (reward asymmetry and burden asymmetry) are introduced and allowed to diffuse through the network. We track the resulting behavioral trajectories to assess whether prosociality is sustained or eroded at the population level. The findings reveal that policy-induced inequities significantly undermine prosocial behavior in LLM agents, an effect amplified through social contagion and mediated by agents’ perceived unfairness.
In summary, our contributions are threefold:
@@
We propose ProSim, a simulation framework for modeling the emergence and evolution of prosocial behavior in LLM agents, integrating four key modules to approximate the complexity of real-world human social environments.
We conduct human benchmarking to validate the capacity of LLM agents to simulate prosocial behavior and to detect and respond to perceived unfairness.
We extend existing theories of prosociality by investigating how structural policy inequities influence the decay and diffusion of prosocial norms within simulated societies.
Related Work
LLM-Driven Agent-Based Social Simulation. Recent progress in large language models (LLMs) has enabled their use as computational agents capable of simulating human-like behaviors across psychology (xu2024ai), economics (horton2023large), and multi-agent systems (guo2024large). LLM-based simulations span three analytical levels. At the individual level, LLMs have been used to model cognitive tasks such as psychological assessments (karra2022estimating), decision-making (horton2023large), and human–computer interactions (farn2023tooltalk; chalamalasetti2023clembench), with growing attention to their metacognitive capacities (zhou2024metacognitive). At the interactional level, LLM agents support studies of coordination (xiong2023examining; qian2023chatdev), moral reasoning (hamilton2023blind; he2024simucourt), and strategic behavior (light2023text; wu2023deciphering). At the societal level, researchers have simulated large-scale dynamics such as norm formation (li2024culturepark), collective action (chuang2023wisdom; zhu2024generative), and opinion diffusion (liu2024skepticism; chuang2023simulating). While these studies demonstrate the broad applicability of LLM agents, their capacity to model diverse prosocial behaviors remains underexplored.
Prosocial Behavior Theory. Prosocial behavior refers to voluntary actions intended to benefit others and plays a key role in fostering social cohesion, cooperation, and trust in institutions (grueneisen2022development). Psychological antecedents include empathic concern (cameron2022empathy; wu2024motive), moral identity (vcehajic2024threaten; yang2025relationship), altruistic orientation (amitha2024altruistic), and perceived social responsibility (pastor2024study; alfirevic2023role). These individual traits interact with contextual factors such as perceived fairness (caserta2023good; tu2022treat), social norms (graf2023social; rudert2022following), and group identity (wang2021socioeconomic; xia2021religious) to shape prosocial intent. Moreover, exposure to antisocial norms or institutional unfairness can significantly suppress cooperative behavior (mekvabishvili2023prosocial; silva2022integrating). While existing studies have established prosociality as a dynamic and socially embedded phenomenon, less is known about how it evolves under policy-induced inequities. Our work contributes to this gap by employing LLM agents to simulate more complex human interactions.
The ProSim Framework
@@
Simulating prosocial behavior poses a fundamental challenge due to its psychological and social complexity. Human prosociality is shaped not only by internal traits such as empathy and moral identity, but also by contextual factors, social influence, and institutional structures. To address this, we introduce ProSim, a simulation framework for studying how prosocial behavior emerges and deteriorates in LLM-based agents across varied social and policy environments. Grounded in social psychology (edelmann2020computational), ProSim supports controlled experimentation at both individual and collective levels. As shown in Figure 1, the framework comprises four modules: (1) Individual Simulation, (2) Scenario Simulation, (3) Interaction Simulation, and (4) Intervention Simulation. These components jointly model the dynamic interplay between psychological traits, situational context, and interventions in shaping agent behavior.
Individual Simulation
@@
This module defines the foundational identity of each agent by capturing the heterogeneity inherent in human populations. Agents are initialized with realistic demographic and psychological profiles:
@@
Psychological Traits. Each agent is further characterized by two sets of psychological traits: (1) core prosocial dispositions (eisenberg1999consistency), including empathic concern, moral identity, altruistic tendency, and social responsibility; and (2) the Big Five personality (gosling2003very), including openness, conscientiousness, extraversion, agreeableness, and neuroticism. Trait values are drawn from Gaussian distributions calibrated using meta-analytic norms to reflect empirical variability.
@@
All attributes are encoded into the agent’s natural language prompt, grounding each agent in a contextually rich identity. This design ensures that behavior arises not from static rules but from socially and psychologically plausible profiles.
Scenario Simulation
@@
This module defines the situational contexts in which agents are prompted to make prosocial decisions. This component is essential for evaluating how different types of social dilemmas and environmental cues influence prosocial tendencies under controlled conditions. To ensure comprehensive coverage of real-world prosocial behaviors, we define six distinct scenarios (Volunteering, Helping, Donating, Cooperation, Recycling, Sharing) that vary across cost level, norm expectation, reward visibility, and dependency level on others. The classification of each scenario along the four dimensions is guided by expert judgment from social psychologists. Each scenario is delivered via a standardized prompt, designed to reflect realistic decision-making contexts. Following the prompt, agents respond using a consistent 7-point Likert scale to indicate their level of prosocial intention.
Interaction Simulation
@@
This module captures how prosocial behavior evolves through repeated agent-to-agent interactions within a structured social network. This layer is essential for modeling emergent dynamics that cannot be explained by individual traits or isolated decisions alone. Concretely, we initialize the agent society using a small-world network generated via the Watts–Strogatz algorithm (kleinberg2000small). This structure reflects key properties of real-world social systems, including high local clustering and short average path lengths. Each node represents an LLM-based agent, and each edge denotes a potential communication or observation channel between two agents.
@@
At each simulation timestep , a random subset of edges is activated. Each agent observes the latest prosocial decisions made by neighbors , where is the set of agents connected to via active edges at time . By integrating (i) the shared scenario narrative , (ii) its own prior decision , and (iii) the observed actions of its neighbors , the LLM agent then generates its decision on six prosocial scenarios: Through this module, we can simulate behaviors dynamics over time within a socially structured community of LLM agents.
Intervention Simulation
@@
This module enables controlled manipulation of policy conditions to investigate how institutional interventions influence the development of prosocial behavior. While individual traits and social dynamics contribute to behavioral variation, policy environments play a central role in shaping collective behaviors through top-down regulation.
@@
Policy inequity. Beyond evaluating idealized interventions, we introduce policy inequity to examine the broader social impact of unfair treatment. In real-world systems, rewards and burdens are often distributed asymmetrically across individuals or groups due to structural bias or implementation constraints. To simulate these conditions, we define two types of inequity: (1)Reward Asymmetry: agents receive unequal recognition or benefits despite contributing equally. (2)Burden Asymmetry: agents face unequal costs or effort for performing the same prosocial task. We randomly assign these asymmetries to a subset of agents and allow their effects to propagate through the social network. This setup enables us to assess how inequitable policies influence the emergence and spread of prosocial behavior within agent societies.
Experimental Settings
@@
To evaluate the capabilities of the ProSim framework, we conduct three progressive studies that examine how LLM agents exhibit, perceive, and respond to prosocial dynamics:
Study 1: Can LLM Agents Exhibit Prosocial Behavior?
Study 2: Do LLM Agents Perceive Inequity?
Study 3: How Does Policy Inequity Affect Prosociality?
@@
Overall Model and Agent Configuration. We evaluate ProSim using five LLMs, including three open-source models (LLaMA-3-8B (llama3), Qwen-2.5-7B (qwen2), DeepSeek-v3 (deepseek)) and two proprietary models (GPT-3.5-turbo (gpt35), GPT-4o (gpt4)), with the generation temperature is set to 0 for reproducibility. We initialize 104 agents sampled from real-world distributions, embedded in a small-world network with neighborhood size and rewiring probability .
@@
Human Benchmarking. We conduct parallel experiments with 104 human participants recruited online for Study 1 and Study 2. Each participant completed the same tasks presented to the LLM agents, using identical scenario prompts and response formats. Prior to task completion, participants provided informed consent and completed standardized psychological inventories aligned with the trait dimensions used in agent initialization. Participants were demographically diverse (Mean age = 32.0 years, SD = 9.4; 53% female), with varying levels of education and occupational backgrounds.
Study 1 Results
Baseline Prosocial Intention in Diverse Scenarios
To establish a foundational comparison between LLM agents and humans, we first assessed baseline prosocial intentions across six representative social scenarios. Each agent responded to a neutral prompt and rated its prosocial behavior on a 7-point Likert scale. Figure 2 (left) presents the distribution of responses for each model across scenarios, while Figure 2 (right) summarizes overall prosociality.
All LLMs demonstrated a general tendency toward prosocial behavior, with average ratings exceeding the midpoint of the scale. Among them, GPT-3.5 exhibited the highest overall prosociality (mean = 4.875), substantially above the human reference (mean = 4.226). Qwen-2.5 (mean = 4.114) and GPT-4o (mean = 4.133) most closely matched the human average, whereas LLaMA-3 (mean = 3.761) and DeepSeek (mean = 3.857) showed lower overall prosocial tendencies. To evaluate not only the magnitude but also the alignment of response patterns, we computed Pearson correlations between each model’s mean ratings and human ratings across the six scenarios. GPT-4o achieved the strongest correlation with human behavior (r = 0.955, p = 0.002), followed by GPT-3.5 (r = 0.923, p = 0.008), and Qwen-2.5 (r = 0.912, p = 0.011). While Qwen-2.5 showed a comparable average score to humans, its higher p-value indicates less reliable alignment across individual scenarios. LLaMA-3 (r = 0.845, p = 0.034) and DeepSeek (r = 0.770, p = 0.072) trailed behind in both magnitude and pattern similarity. These results suggest that state-of-the-art LLMs are capable of expressing prosocial intentions that not only approximate human averages but also exhibit scenario-level consistency with human judgments.
The Influence of Psychological Traits on Prosociality
@@
Behavioral Shifts Under Policy Interventions
We next evaluate whether LLM agents adjust their prosocial behavior when exposed to external policies. Each agent completes the same six scenarios from the baseline condition, with prompts modified to include one of four intervention framings. We compute the relative change in prosocial intention to assess policy responsiveness. Figure 4 summarizes the results. Intervention effectiveness varies across policy types and models. Regulatory Enforcement yields the most consistent improvement across models and scenarios, suggesting that explicit institutional mandates are broadly effective at eliciting compliance. In contrast, Economic Incentives, while highly effective for human participants, produce mixed effects in LLMs. Only GPT-4o and DeepSeek-v3 exhibit reliable prosocial gains across all six scenarios, though the magnitude of their responses remains below human levels, indicating that monetary cues are less salient for LLMs. The right panel of Figure 4 aggregates the number of positive shifts across all policy-scenario combinations, also showing that these two models can exhibit the strongest responsiveness, closely matching human participants. Taken together, these findings suggest that advanced LLMs can exhibit human-like adjustments to social and normative inputs.
Study 2 Results
This section examines whether LLM agents are capable of recognizing unfairness and responding in norm-enforcing ways. To evaluate this capacity, we adapt a third-party punishment paradigm from behavioral economics (fehr2004third), a well-established framework for studying fairness judgment and altruistic punishment. Each agent completes a 60-trial task in which two newly assigned virtual players propose an allocation of $30. Player 1 offers $x to Player 2 and retains $[30-x] for themselves. The participants, acting as third-party judges, then chose between:
@@
We record agents’ binary choices (accept = 0, punish = 1) across trials and and compare the results with humans. Figure 5 illustrates punishment rates across different combinations of fairness levels and punishment costs for each model. Human participants exhibit a clear normative pattern: punishment rates decrease as allocations become more equitable or as costs increase. This behavior reflects a trade-off between norm enforcement and self-interest, consistent with findings from economic game theory. Among LLMs, the GPT series shows the strongest human alignment. These models reliably punish under unfair conditions and abstain under fair ones, demonstrating a consistent threshold-based fairness judgment. Notably, they even exceed human punishment rates in highly unfair trials. DeepSeek-v3 maintains high punishment rates across all unfair conditions, largely ignoring cost variation, suggesting a rigid but robust norm-enforcing strategy. LLaMA-3 also punishes inequity frequently, but its responses are less differentiated by fairness level or cost. Qwen-2.5, meanwhile, shows low overall punishment and lacks sensitivity to either factor. These findings indicate that while some advanced LLMs demonstrate fairness-based reasoning and partial cost sensitivity, others lack the granularity and flexibility observed in human social decision-making.
Study 3 Results
Impact of Inequity on Prosocial Behavior